# custom.bib

### v1

@@ 5.1 Accuracy
For 2D images, we show that the LieRE-based transformer outperforms the DeiT by 5.5% [touvron2022deit], RoPE adaptation in VisionLlama [chu2024visionllama] by 3.6% and RoPE-Mixed by 2.7% on Cifar100 [heo2024rotary].
For 3D inputs (video and 3D medical images), we observe an accuracy improvement of LieRE-based transformer over the Absolute Position Embedding of up to 6.7 % and RoPE-based transformers by up to 2.5% in accuracy.
Table 1: 2D image classification on Cifar100 and ImageNet [krizhevsky2009learning, deng2009imagenet], equivalent to DeiT
Method # params Cifar100 Accuracy (%) ImageNet Accuracy (%)
Absolute Position E. 85.1M 63.9 66.1
VisionLlama RoPE 85.1M 65.8 -
RoPE-Mixed 85.1M 66.7 68.5
LieRE (ours) 85.1M 69.4 68.8
Table 2: 3D video classification on RSNA hemorrhage prediction and UCF101 dataset [soomro2012ucf101, rsna2019data], equivalent to Vivit (spatio-temporal)
Method # params UCF101 Accuracy (%) RSNA Accuracy (%)
Absolute Position E. 88.7M 44.4 80.7
RoPE-Mixed 88.7M 48.6 81.9
LieRE (ours) 88.7M 51.1 82.7
@@ 5.2 Data efficiency
We observe that transformers based on LieRE exhibit greater data efficiency compared to leading transformer architectures for 2D images on the CIFAR-100 dataset. For instance, using only 40% of the dataset, the LieRE-based transformer outperforms all comparable methods by 4.47%. Additionally, there is a mild trend indicating that the performance gap widens as the amount of data decreases (Figure 2(b)).
(a) Performance scales with generator dimension (learnable parameters)
(b) Data ablation on Cifar100
Figure 2:
5.3 Generator Dimension Scaling Laws
LieRE adds a small amount of capacity to the model ($4032$ parameters in our case) but still represents the largest amount of learnable capacity around relative position encodings to date. This leads to the natural question of how helpful marginal capacity is.
We vary capacity by enforcing a block diagonal structure on the generator. varying the block size allows us to approach "full" LieRE. Note that $2\times 2$ blocks correspond exactly to RoPE-Mixed.
@@ 5.4 Compute efficiency
Figure 3: The LieRE spatial encoding allows the model to match the performance of absolute position encodings with substantially less training time.
Training transformers can necessitate substantial computational resources, which can hinder equitable access to research and development of machine learning methods. We demonstrate that the LieRE-based transformer requires 3.5 times less training time to achieve comparable performance to the Absolute Position Embedding baseline (as used in DeiT III [touvron2022deit]). This represents the largest reduction in training time compared to recent works such as VisionLlama and RoPE-Mixed [chu2024visionllama, heo2024rotary]. We note that this is the first training efficiency comparison of these recent methods. Figure 3 shows the amount of allowable compute reduction to achieve the same accuracy achieved by absolute position encodings (DEIT baseline) after 200 epochs. LieRE demonstrates the largest win, allowing a 3.5X reduction in training compute to achieve the same accuracy.
While all methods display better performance with more training compute, a $3.5x$ reduction allows for significantly faster experimentation and makes a fixed level of performance more accessible to those with limited access to compute resources.
@@ 5.5 Patch shuffling
We evaluate models using the decline in accuracy when evaluating on shuffled patches. We observe the most significant decline LieRE-based transformers, leading to the conclusion that LieRE encodes meaningful relative positions between tokens. The complete results are displayed in table LABEL:tab:shuffling_results for cifar100 and table LABEL:tab:shuffling_results_3D for UCF101.
Table 3: Relative drop in accuracy for 2D image classification after shuffling patch order, Cifar100
Method before shuffling $\uparrow$ after shuffling $\downarrow$ Accuracy Drop (%) $\uparrow$
Absolute Position E. 63.9 19.6 69.3
VisionLlama RoPE 65.8 29.7 54.8
RoPE-Mixed 66.7 16.5 75.4
LieRE (ours) 69.4 15.9 77.1
Table 4: Accuracy for Video recognition after patch shuffling, UCF101 dataset
Method before shuffling $\uparrow$ after shuffling $\downarrow$ Accuracy Drop (%) $\uparrow$
Absolute Position E. 44.4 44.0 0.1
RoPE-Mixed 48.6 36.6 24.7
LieRE (ours) 51.1 32.2 36.9
@@ 6 Limitations
While LieRE shows promising results across multiple modalities and input dimensionalities, there are a few limitations worth noting:
1. Compatibility with architectures that do not use attention: Our method is specifically designed to modify the inner product, making it compatible with most attention schemes, including original attention and linear attention. However, this specificity may limit its applicability to other architectures, like convolutional neural networks, that do not rely on inner product-based attention mechanisms. Future work could explore adaptations of our method to a wider range of architectures.
2. Limited to positions in $\mathbb{R}^{n}$: The current formulation of our method is designed to encode positions in $\mathbb{R}^{n}$. While this is sufficient for many applications, it may not be directly applicable to tasks that require encoding poses in $SE(3)$, like robotics. Further research may be necessary to adapt our method to effectively handle such representational requirements.
@@ 7 Broader Impacts
LieRE, our proposed method for encoding positional information in attention mechanisms, has demonstrated substantial improvements in 2D and 3D image classification tasks. We are particularly excited about how LieRE can expand the applicability of RoPE-based transformers to $n$ dimensions. This lays down the tracks to apply that same relative position within and across modalities and settings.
1. One Backbone for All Modalities: LieRE enhances the generalizability of attention mechanisms, making them more versatile and adaptable to a wider range of modalities. Once the input is encoded as tokens and positions, no additional changes are necessary in order to support a new modality. By providing a flexible and efficient way to encode positional information, LieRE can facilitate the application of attention mechanisms to diverse data types, such as audio, video, multi-dimensional sensor data or temporal data with diverse time intervals. This increased generalizability can lead to more effective transfer learning and cross-pollination of ideas between different research areas, accelerating progress and innovation in the field.
2. Improving accessibility of attention-based architectures for low-data and low-compute regimes: This is particularly valuable for applications where labeled spatial data is scarce or expensive to acquire, such as in medical imaging and robotics. By lowering the barriers to entry, LieRE can enable the application of attention mechanisms to a broader set of problems, including those in resource-constrained settings or niche domains.
3. Enabling more efficient and sustainable AI: The reduced training costs and improved data efficiency associated with LieRE can contribute to the development of more efficient and sustainable AI systems. By requiring fewer computational resources and less training data, LieRE can help mitigate the environmental impact of large-scale model training. This is particularly important as the field of AI continues to grow and the demand for computational resources increases.
In conclusion, LieRE has the potential to expand the generalizability of attention mechanisms, improve their accessibility for spatial data, and contribute to the development of more efficient and sustainable AI systems.
@@ 8 Conclusion
In this paper, we introduced Lie group Relative position Encodings (LieRE), a position encoding that can effectively encode relative position information for attention mechanisms across modalities and input dimensionalities. Through extensive experiments on 2D image classification (CIFAR100, ImageNet) and 3D medical image classification (RSNA brain CT hemorrhage detection), we demonstrated that LieRE achieves superior performance compared to existing positional encoding methods. Beyond improving accuracy, LieRE also exhibits data and compute efficiency. On CIFAR100, LieRE requires 3.5 times less training computed to match the performance of the baseline model with absolute position encodings. Furthermore, LieRE can outperform the baseline trained on the full dataset while using only 70% of the training data, highlighting its data efficiency. The key advantages of LieRE include its simplicity, flexibility, and ease of adaptation to new modalities. By requiring no changes to the tokenizer other than outputting positions in $\mathbb{R}^{n}$ and no other code changes, LieRE provides a unified and efficient approach for transformers to process and learn from various data modalities within a single architecture.
We hope that this empowers the field in building models that can understand concepts.
Funding and Acknowledgments We would like to thank Maksim Maydanskiy for helping us understand Lie groups and Lie group representations. We would like to thank Aradhana Sinha for suggesting the shuffling experiment and providing early feedback on the paper.
Sophie Ostmeier was supported by the German Research Foundation (DFG), Walter-Benjamin fellowship (ID: 517316550).
This project was supported by Google Cloud and Azure credits.
@@
Appendix A Experiment Hyperparameters
The backbone for all experiments is configured as ViT-B, with 12 layers, a hidden dimension of 768, and an intermediate dimension of 3096. We use dropout of 0.1.
A.1 2D Image Classification
The CIFAR experiments where trained on 8xL4 GPUs with 24GB of VRAM each and all took under 30 minutes to complete. The generator scaling experiment was conducted using RTX6000 GPUs. The ImageNet experiments were trained on 8xL40 GPUs and all took less than 2 days and 5 hours of runtime including time lost due to preemption and resource sharing. We use a cosine learning rate schedule with an initial learning rate of $41E-$ 4 and train for 200 epochs. We use an effective batch size of $512$. We use a patch size of $4\times 4$ on the original $32\times 32$ image for Cifar100 and a patch size of $16\times 16$ on the randomly cropped and resized $224\times 224$ image. All vision experiments used RandAugment [cubuk2020randaugment]. We use the ADAM optimizer with betas of 0.9 and 0.999 and $\epsilon=1e-8$. The hyperparameters were tuned with RoPE-mixed and selected before conducting the LieRE trainers as to ensure a fair comparison.
A.2 3D Medical Image and Video Classifications
The 3D classification experiments were conducted on either $1008\times A$ 100 40GB GPUs or $1004\times A$ 100 80GB GPUs with the effective batch size held constant either by using a gradient accumulation or increasing the batch size. Similar to 2D classification, we use an initial learning rate of $41E-$ 4 with a cosine decay, trained for $200$ epochs, and had a total batch size of $64$ and a patch size of $4\times 16\times 16$ on the randomly cropped and resized $32\times 224\times 224$ video/image. We use the ADAM optimizer with betas of 0.9 and 0.999 and $\epsilon=1e-8$. We use Mixup for all 3D experiments [tong2022videomae, zhang2017mixup].

### v2

@@ 5.1 Accuracy
For 2D images, we show that the LieRE-based transformer outperforms the DeiT by a relative performance improvement of 9.4% [touvron2022deit], RoPE adaptation in VisionLlama [chu2024visionllama] by 6.7% and RoPE-Mixed by 9.7% on CIFAR-100 [heo2024rotary].
For 3D inputs (video), we observe an relative accuracy improvement of LieRE-based transformer over the absolute position embedding of up to 25.5 % and RoPE-inspired position encodings at least to 6.1% in accuracy.
Table 1: 2D image and 3D video classification Top-1 Accuracy (95% confidence intervals) results. All models use 85.1M parameters for 2D tasks and 88.7M parameters for 3D task [krizhevsky2009learning, deng2009imagenet, soomro2012ucf101, rsna2019data] equivalent to DeiT, ∗∗ equivalent to Vivit (spatio-temporal).
Method CIFAR-100 ImageNet-1k UCF101
Abs. Pos. E.∗,∗∗ 63.9 (62.9-65.8) 66.1 (65.7-66.5) 42.7 (42.3-43.1)
VisionLlama RoPE 65.5 (64.6-66.5) 65.4 (65.0-65.8) 50.5 (50.1-50.9)
RoPE-Mixed 63.7 (61.8-65.6) 65.9 (65.5-66.3) 45.7 (45.2-46.0)
LieRE-Commute 66.9 (66.0 - 67.9) 68.4 (67.9-68.9) 52.1 (51.7-52.5)
LieRE 69.9 (68.9-70.8) 69.5 (69.1-69.9) 53.6 (53.2-54.0)
@@ 5.2 Data efficiency
We observe that transformers based on LieRE exhibit greater data efficiency compared to leading transformer architectures for 2D images on the CIFAR-100 dataset. For instance, when using 40% of the dataset, the LieRE-based transformer outperforms all comparable methods by 4.5 percentage points. Additionally, there is a mild trend indicating that the performance gap widens as the amount of data decreases (Figure 3(b)).
(a) Performance scales with generator dimension the number of LieRE parameters, compared to LieRE with one skew-symmetric matrix learned for the entire model (stem).
(b) Data ablation on for different position embeddings on CIFAR-100.
5.3 Model Scaling
We examine three axes of capacity scaling: transformer backbone parameters, LieRE generator parameters and learning separate LieRE generators across heads and layers. Here we find a surprising result that RoPE-mixed can be substantially improved by sharing parameters across attention heads and layers and not learning them separately.
5.3.1 Transformer backbone capacity
We also evaluate how the inclusion of LieRE affects performance across model sizes on the CIFAR-100 in Figure 4(a). We observe that LieRE retains a statistically significant lead in performance across all three model sizes.
Refer to caption (a) Performance behavior on CIFAR-100 (2D Image Classification) over ViT-Tiny (22M), ViT-Base (85M), ViT-Large (302M) for LieRE Rope-Mixed and Absolute Encoding (Appendix, table 5).
Refer to caption (b) The LieRE spatial encoding allows the model to match the performance of absolute position encodings with substantially less training time.
5.3.2 Generator parameters
LieRE adds a small amount of capacity to the model ($4032$ parameters for the ViT-B backbone we use for most experiments), leading to the natural question of how helpful the marginal capacity is.
We vary capacity by enforcing a block diagonal structure on the generator. varying the block size allows us to approach "full" LieRE. Recall that using $2\times 2$ blocks recovers exact commutativity and is referred to as LieRE-commute. This configuration substantially outperforms RoPE-Mixed, despite differing only in that it uses the same parameters across heads and layers and RoPE-Mixed learns separate parameters.
@@ 5.3.2 Generator parameters
5.3.3 Sharing LieRE Parameters Across Heads and Layers
We also examine whether learning separate positional encoding per attention head and per layer improves performance on CIFAR-100 in table 2. We conduct this experiment for both RoPE-Mixed and LieRE and observe that in both cases it is best to learn one set of parameters across heads and layers. This is the mechanism by which we are able to show larger improvements from adopting RoPE-Mixed than the original paper [heo2024rotary]. We name this transformed form of RoPE-Mixed LieRE-Commute, since it is equivalent to LieRE with generator dimension 2 (see Figure LABEL:fig:generator_dataset).
Table 2: Layer, Head and Stem Learnable Rotary Embeddings on CIFAR-100, ViT-B, equivalent to LieRE-Commute
FLOP Stem Heads Layers Rope-Mixed LieRE Reference
5.684G 63.0 64.5 [heo2024rotary]
5.684G 63.6 65.1
5.613G 68.3 69.9
5.613G 66.9 69.9 (ours)
@@ 5.4 Compute efficiency
Training transformers can necessitate substantial computational resources, which can hinder equitable access to research and development of machine learning methods. We demonstrate that the LieRE-based transformer requires 3.5 times less training epochs on Cifar100 to achieve comparable performance to the Absolute Position Embedding baseline (as used in DeiT III [touvron2022deit]). This represents the largest reduction in training time compared to recent works such as VisionLlama and RoPE-Mixed [chu2024visionllama, heo2024rotary]. We note that this is the first training efficiency comparison of these recent methods. Figure 4(b) shows the amount of allowable compute reduction to achieve the same accuracy achieved by absolute position encodings (DeiT baseline) after 200 epochs. LieRE demonstrates the largest win, allowing a 3.5X reduction in training compute to achieve the same accuracy.
5.5 FLOP analysis
We find that since all methods we examine introduce a computational cost that is at most linear in the number of tokens, and runtime is dominated by the quadratic attention component, there is no substantial difference in computational efficiency between the methods. We list inference FLOP of the various methods in table 3.
Table 3: FLOP analysis with percentage increase compared to absolute position encodings
Position Enc. ViT-Tiny (22M) ViT-Base (85M) ViT-Large (302M)
Abs. Pos. E. 0.963G 5.607G 19.856G
VisionLlaMA RoPE 0.963G (+0.001%) 5.607G (+0.002%) 19.856G (+0.000%)
RoPE-Mixed 0.963G (+0.001%) 5.609G (+0.036%) 19.863G (+0.035%)
LieRE-Commute (ours) 0.963G (+0.000%) 5.608G (+0.018%) 19.858G (+0.005%)
LieRE (ours) 0.970G (+0.727%) 5.613G (+0.107%) 19.869G (+0.065%)
@@ 5.6 Patch shuffling
We evaluate models using the decline in accuracy when evaluating on shuffled patches. We observe the most significant decline LieRE-based transformers, leading to the conclusion that LieRE models are most capable at using positional information. The complete results are displayed for CIFAR-100 and table for UCF101 (Table 4) .
Table 4: Relative accuracy drop for 2D image classification (CIFAR-100) and Video recognition (UCF101) after patch shuffling
CIFAR-100 (2D) UCF101 (3D)
Method Before After Drop(%) Before After Drop(%)
[…]
Abs. Pos. E. 63.9 19.6 69.3 42.7 42.5 0.0
VisionLlama RoPE 65.8 29.7 54.8 50.5 36.2 28.5
RoPE-Mixed 63.6 38.7 39.0 45.7 37.9 17.1
LieRE-Commute 66.9 16.5 75.4 52.1 37.2 28.6
LieRE 69.9 15.9 77.1 53.6 24.7 53.9
@@ 6 Limitations
While LieRE shows promising results across multiple modalities and input dimensionalities, there are a few limitations worth noting. Our method is specifically designed to modify the inner product, making it compatible with most attention schemes, including original attention and linear attention. However, this specificity may limit its applicability to other architectures, like convolutional neural networks, that do not rely on inner product-based attention mechanisms. Future work could explore adaptations of our method to a wider range of architectures.Furthermore, the current formulation of our method is designed to encode vector positions in $\mathbb{R}^{d}$. While this is sufficient for many applications, it may not be directly applicable to tasks that require encoding poses in $SE(3)$ like robotics. Further research may be necessary to adapt our method to effectively handle such representational requirements.
@@ 7 Broader Impacts
LieRE, our proposed method for encoding positional information in attention mechanisms, has demonstrated improvements in 2D image and video classification tasks. We are particularly excited about how LieRE can expand the applicability transformers to $n$ dimensional inputs. This may lay down the tracks to apply that same relative position encoding within and across modalities and settings.
@@ 8 Conclusion
In this paper, we introduced Lie group Relative position Encodings (LieRE), a position encoding that can effectively encode relative position information for attention mechanisms across modalities and input dimensionalities. Through experiments on 2D image classification (CIFAR-100, ImageNet-1k) and 3D video classification (UCF101), we demonstrated that LieRE achieves better performance compared to existing positional encoding methods. Beyond improving accuracy, LieRE also exhibits data and compute efficiency. On CIFAR-100, LieRE requires 3.5 times less training computed to match the performance of the baseline model with absolute position encodings. Furthermore, LieRE can outperform the baseline trained on the full dataset while using only 70% of the training data, highlighting its data efficiency. The key advantages of LieRE include its simplicity, flexibility, and ease of adaptation to new modalities. By requiring no changes to the tokenizer other than outputting positions and no other code changes, LieRE may provide a unified and efficient approach for transformers to process and learn from various data modalities within a single architecture.
@@
Appendix A Appendix
Appendix B Experiment Hyperparameters
The backbone for all experiments is configured as ViT-B, with 12 layers, a hidden dimension of 768, and an intermediate dimension of 3096. We use a dropout of 0.1. We used CLS pooling in our implementation to facilitate comparability with existing literature in the field. Further experiments revealed substantial performance improvement with mean pooling and LieRE. We use the pytorch lightning framework for all experiments [falcon2019pytorch].
B.1 2D Image Classification
The CIFAR experiments where trained on 8xL4 GPUs with 24GB of VRAM each and all took under 30 minutes to complete. The generator scaling experiment was conducted using RTX6000 GPUs. The ImageNet experiments were trained on 8xL40 GPUs and all took less than 2 days and 5 hours of runtime including time lost due to preemption and resource sharing. We use a cosine learning rate schedule with an initial learning rate of $41E-$ 4 and train for 200 epochs. We use an effective batch size of $512$. We use a patch size of $4\times 4$ on the original $32\times 32$ image for CIFAR-100 and a patch size of $16\times 16$ on the randomly cropped and resized $224\times 224$ image. All vision experiments used RandAugment [cubuk2020randaugment]. We use the ADAM optimizer with betas of 0.9 and 0.999 and $\epsilon=1e-8$. The hyperparameters were tuned with RoPE-mixed and selected before conducting the LieRE trainers as to ensure a fair comparison.
B.2 3D Video Classifications
The 3D classification experiments were conducted on either $1008\times A$ 100 40GB GPUs or $1004\times A$ 100 80GB GPUs with the effective batch size held constant either by using a gradient accumulation or increasing the batch size. Similar to 2D classification, we use an initial learning rate of $41E-$ 4 with a cosine decay, trained for $200$ epochs, and had a total batch size of $64$ and a patch size of $2\times 16\times 16$ on the randomly cropped and resized $32\times 224\times 224$ video/image. We use the ADAM optimizer with betas of 0.9 and 0.999 and $\epsilon=1e-8$.
B.3 Scaling Model Size
Table 5: Comparison of Position Encoding Methods for Different ViT Models, Accuracy (bootstrapped 95%CI)
Position Encoding ViT-Tiny ViT-Base ViT-Large
Abs. Pos. E. 57.2 (56.2 - 58.1) 63.9 (62.9 - 65.8) 60.5 (59.5 - 61.4)
VisionLlaMA RoPE 58.2 (57.2 - 59.2) 65.5 (64.6 - 66.5) 62.3 (60.4 - 64.2)
RoPE-Mixed 59.6 (58.7 - 60.6) 63.7 (61.8 - 65.6) 65.0 (64.0 - 65.9)
LieRE-Commute 63.2 (62.2 - 64.1) 66.9 (66.0 - 67.9) 65.1 (64.1 - 66.0)
LieRE 64.7 (63.8 - 65.7) 69.9 (68.9 - 70.8) 69.2 (68.3 - 70.1)
B.4 Python implementation of skew-symmetric matrix
1generator_raw_params = nn.Parameter(
2 torch.rand(
3 input_dimensionality,
4 head_dim,
5 head_dim,
6 ) * 2 * math.pi # optional, inspired from RoPE-Mixed paper
7)
8
9upper_triangle = (
10 torch.triu(generator_raw_params, diagonal=1)
11)
12skew_bases = upper_triangle - torch.transpose(upper_triangle, -1, -2)
13in_basis_positions = (
14 positions.reshape(list(positions.shape) + [1] * 2) * skew_bases
15)
16generator_pos = torch.sum(in_basis_positions, dim=-3)
17rotation = torch.matrix_exp(generator_pos.to(dtype=torch.float32)).to(dtype=positions.dtype)